# Cost-aware LLM-based Online Dataset Annotation

### v1

@@ E.1 Additional Experimental Results
Figure 11: Cumulative average accuracy (blue) and cost (red) of CaMVo with $\delta=0.96$, $k_{\min}=1$ under nine different permutations of the dataset. The green line marks each $\delta$-specific target accuracy.
@@
Figure 12: Cumulative average accuracy (blue) and cost (red) of CaMVo with $\delta=0.96$, $k_{\min}=3$ under nine different permutations of the dataset. The green line marks each $\delta$-specific target accuracy.
Figure 8 visualizes CaMVo’s learning trajectories on IMDB for $k_{\min}=1$ across all the confidence thresholds $\delta$ values reported in Table 4. In all cases, CaMVo begins by querying larger, higher-cost ensembles to obtain reliable performance estimates, then rapidly shifts to cost-optimal subsets once the lower-confidence bounds converge. This transition yields a sharp decline in cost while maintaining accuracy above the target line.
At the extreme threshold $\delta=0.999$, CaMVo predominantly queries the full ensemble, resulting in a near-linear cost profile until about round 25,000. For $\delta\leq 0.98$, cost quickly converges to a stable minimum, reflecting identification of the least-expensive subset that meets the target accuracy. Across all plots, CaMVo achieves or exceeds the respective accuracy target. For very high thresholds ($\delta\geq 0.995$), final accuracy hovers just above the threshold, as expected; as $\delta$ decreases, the accuracy surplus grows. Below $\delta=0.96$, accuracy plateaus at approximately 94.06%, corresponding to the performance of the single cheapest model (‘llama-3.1-8b’).
Overall, these results demonstrate CaMVo’s ability to balance exploration and exploitation, swiftly discover cost-effective subsets, and reliably satisfy the target accuracy requirements.
Figure 10 presents CaMVo’s learning curves under varying confidence thresholds $\delta$ with $k_{\min}=3$. Similar to the $k_{\min}=1$ case, CaMVo initially selects larger, more expensive ensembles to obtain reliable performance estimates, then quickly transitions to lower-cost subsets as the lower confidence bounds stabilize. For $\delta\geq 0.99$, the cost and accuracy trajectories are nearly identical to those observed with $k_{\min}=1$. When $\delta\leq 0.985$, CaMVo settles on the lowest-cost subset of three LLMs, and the plots when $\delta\leq 0.985$ are almost-identical.
@@
To assess CaMVo’s sensitivity to dataset ordering, Figure 10 plots the mean cumulative accuracy and cost curves (solid lines) for MMLU when $\delta=0.995$ , $k_{\min}=1$ (Left); and $\delta=0.995$ , $k_{\min}=1$ (Right), averaged over 20 random shuffles. Shaded regions indicate one standard deviation. Although the accuracy band starts wide, which reflects both the initial exploration and also the varying mixes of ’easier’ and ’harder’ instances; this rapidly contracts over time, confirming CaMVo’s reliable attainment of the target accuracy across permutations. The cost band likewise narrows, demonstrating stable convergence to low-cost ensembles. Notably, the accuracy variability remains much smaller than the cost variability, as CaMVo does not optimize toward a fixed cost. Comparing $k_{\min}=1$ and $k_{\min}=3$, it can be seen that both the cost and accuracy bands are narrower, as there are less available subsets to choose when $k_{\min}=3$.

### v2

@@ Results.
Figure 10: Cumulative average accuracy (blue) and cost (red) of CaMVo ($k_{\min}=1$) on the AG News Classification Dataset across rounds for various confidence thresholds $\delta$. The green line marks each $\delta$-specific target accuracy.
Figure 10 visualizes CaMVo’s learning trajectories on the AG News Classification Dataset for $k_{\min}=1$ across all the confidence thresholds $\delta$ values reported in Table 4. Similar to experiments on the other datasets, CaMVo begins by querying larger, higher‐cost ensembles to obtain reliable performance estimates, then rapidly shifts to cost-optimal subsets once the lower‐confidence bounds converge.
Figure 11: Mean (solid lines) and one‐standard‐deviation bands (shading) of CaMVo’s cumulative average accuracy (blue) and cost (red) over 20 random shuffles using the AG News Classification Dataset ($\delta=0.99$, $k_{\min}=1$). The green line indicates the accuracy target of $95.14\%$ for $\delta=0.995$.
@@
Figure 12: Cumulative average accuracy (blue) and cost (red) of CaMVo with $\delta=0.99$, $k_{\min}=1$ under nine different random permutations of the dataset. The green line marks each $\delta$-specific target accuracy.
@@
Further, to assess CaMVo’s sensitivity to dataset ordering, Figure 11 plots the mean cumulative accuracy and cost curves (solid lines) for $\delta=0.99$, $k_{\min}=1$, averaged over 20 random shuffles. Shaded regions indicate one standard deviation. Similar to the experiments in other datasets, although the accuracy band starts wide, which reflects both the initial exploration and also the varying mixes of ’easier’ and ’harder’ instances; this contracts over time, confirming CaMVo’s reliable attainment of the target accuracy across permutations. The cost band likewise narrows, demonstrating stable convergence to low‐cost ensembles.